A gold mine extract gold from rock using toxic substances. For each additional tonne of rock the marginal benefit of extraction are MB =12-q where q is the quantity of rock. The extraction process causes marginal environmental damage equal to MD = 2 q. a) Without regulation, how much rock would the gold mine extract? b) With strict liability laws, how much gold would the mine extract?
Answer #1

a) without regulation

MB = 12- q

MC= 0

therefore gold will be extracted until MB=MC

12-q=0

q= 12

without regulation 12 units of rock will be extracted by the gold mine.

....

b) With strict liability laws

MB= 12-q

MC= 2q

therefore gold will be extracted until MB=MC

12-q = 2q

3q = 12

q = 4

With strict liability laws, 4 units of rock will be extracted by the gold mine.

.....